OpenStudy (mollynicole2015):

Solve for g. g – 3 = 2g

OpenStudy (anonymous):

1=g

OpenStudy (gabebae):

You're not supposed to give out direct answers @xav please expain

OpenStudy (anonymous):

ok then u explain it plus it was an hour ago like calm down man

OpenStudy (anonymous):

g-3=2g -3=2g+g -3=3g -1=g

OpenStudy (anonymous):

Move g to the RHS of the equation. -3 = 2g -g g = -3

OpenStudy (anonymous):

@ninamae05 When you moved -3 to the RHS, you forgot to change it's sign.

OpenStudy (mollynicole2015):

@RobTobey Thank you. That helps. ;)

OpenStudy (anonymous):

Your welcome.